Friedmania is a country in which the quantity theory of money operates. The country has a constant population, capital stock, and technology so real GDP does not change. In 2010, real GDP was $500 million, the price level, measured by the GDP deflator, was 150 and the velocity of circulation of money was 10. (Because the price level is measured by the GDP deflator, it must be divided by 100 before it is used in the equation of exchange.) In 2011, the quantity of money increased by 20 percent.

a) What was the quantity of money in 2010?

b) What was the velocity of circulation in 2011?

c) What was the price level in 2010?


1
Expert's answer
2021-09-12T16:30:05-0400

(a) "Money supply= \frac{Price Level\times RealGDP}{Velocity of Money}"

"= \frac{1.5\times500}{10}=75"

(b)

"Money Supply = \frac{75\times120}{100}=90"

Velocity of cicurlation "= \frac {1.5\times500}{90}=8.3"


(c) Price level in 2010

"= \frac{150}{100}= 1.5"